# BrowseComp-V3: A Visual, Vertical, and Verifiable Benchmark for Multimodal Browsing Agents

### v1

@@ 1 Introduction
Multimodal large language models OpenAI [2025b]; Pichai et al. [2025]; Bai et al. [2025]; Meta [2025]; Li et al. [2024a]; Shi et al. [2025b] have demonstrated substantial performance gains across complex tasks. By integrating linguistic comprehension, visual perception, and tool-use capabilities, these models are increasingly evolving into autonomous agents capable of independent exploration and decision-making. Consequently, an increasing body of research is exploring how MLLMs can leverage external search and browsing tools to address multimodal deepsearch challenges in open-world environments OpenAI [2025a]; Google [2024]; Wu et al. [2025b]; Geng et al. [2025]; Huang et al. [2026].
Despite the rapid evolution of model capabilities, benchmarks for multimodal browsing and deep search remain noticeably underdeveloped. Existing studies Jiang et al. [2024]; Geng et al. [2025]; Li et al. [2025b]; Tao et al. [2025a] frequently exhibit shortcomings in task complexity, information searchability, and evaluation dimensions, hindering fair, holistic, and reproducible assessments of multimodal browsing agents. Existing methods still exhibit certain limitations: i) Insufficient Task Complexity. Early benchmarks Cheng et al. [2025]; Geng et al. [2025] are predominantly confined to shallow retrieval within two hops, with visual information concentrated in the initial stage. Consequently, they fail to reflect the intricacies of real-world, deep multimodal search scenarios. ii) Inaccessibility of Key Information. The core evidence in subsequent benchmarks Fu et al. [2025b]; Li et al. [2025b]; Tao et al. [2025a] is often derived from sources that are not publicly searchable by tools, such as videos or proprietary documents, which undermines the reproducibility and fairness. iii) Narrow Evaluation Dimensions. Existing studies Jiang et al. [2024]; Li et al. [2024b] primarily focus on the accuracy of the final answer but lack a systematic characterization of the reasoning process. This makes it challenging to diagnose specific failure modes or define the capability boundaries of the models.
@@ 2.1 Multimodal Large Language Models
Multimodal large language models OpenAI [2025b]; Pichai et al. [2025]; ByteDance Seed [2025]; Bai et al. [2025] have demonstrated remarkable proficiency across a diverse spectrum of tasks, such as VQA Fu et al. [2025a]; Cheng et al. [2025]; Zhang et al. [2025]; Wu et al. [2025b], grounding Kazemzadeh et al. [2014], OCR Masry et al. [2022]; Shi et al. [2025a]; Mathew et al. [2021]; Shi et al. [2025c], and multimodal reasoning Lu et al. [2023]; Wang et al. [2025, 2024]. Nevertheless, MLLMs inherently struggle with the real-time acquisition of up-to-date information, posing substantial hurdles when addressing knowledge-intensive or information-retrieval queries. Consequently, contemporary research has pivoted toward tool-augmented frameworks to empower MLLMs as autonomous agents, capable of dynamically retrieving and incorporating external knowledge.
@@ 2.2 Tool-Enhanced Browsing Agents
Driven by the escalating tool-calling proficiency of LLMs/MLLMs Guo et al. [2025]; Yang et al. [2025]; OpenAI [2025b], tool-enhanced browsing agents have emerged as a pivotal research frontier. To enable precise retrieval and reasoning in dynamic web environments, recent studies advocate leveraging supervised fine-tuning and reinforcement learning to enhance agents’ reasoning and decision-making capabilities Jin et al. [2025]; Li et al. [2025a]; Wu et al. [2025a]; Tao et al. [2025b]; Song et al. [2025]; Zheng et al. [2025]. This paradigm, initially validated in textual agents, has been rapidly extended to the multimodal domain Wu et al. [2025b]; Mei et al. [2025]; Hong et al. [2025]; Huang et al. [2026]. This evolution has significantly expanded the search depth and adaptive boundaries of agents when navigating complex tasks.
@@ 2.3 Multimodal Browsing Benchmarks
Traditional multimodal browsing benchmarks typically decouple visual understanding from text retrieval and focus on simple two-hop retrieval tasks Chen et al. [2023]; Mensink et al. [2023]; Cheng et al. [2025]; Geng et al. [2025]. As visual agents advance, performance on such tasks has largely saturated. BrowseComp Wei et al. [2025] evaluates text-only agents in open-world settings by requiring large-scale web navigation, offering valuable guidance for multimodal task construction. Inspired by this paradigm, recent benchmarks such as MM-BrowseComp Li et al. [2025b] and MMSearch-Plus Tao et al. [2025a] incorporate multi-hop designs and fine-grained visual reasoning to enhance reasoning depth. However, existing benchmarks still suffer from key limitations: critical information often resides in videos or non-searchable documents, tool support is insufficient, and evaluation primarily measures final-answer correctness while overlooking the quality of reasoning. To bridge this gap, we propose BrowseComp-V3, ensuring all critical information comes from publicly accessible resources during task design. We also introduce the Process Score metric to evaluate multimodal browsing agents comprehensively.
@@ 4.1 Experimental Setup
  • Tool-Free MLLMs. We benchmark multiple SOTA MLLMs in a tool-free setting, where models must generate answers directly without access to external tools or search capabilities. Specifically, we evaluate the following models: GPT-5.2 OpenAI [2025b], o4-mini OpenAI [2025c], GPT-4o Hurst et al. [2024], Gemini-3-Flash-Preview Google [2025], Claude-Sonnet-4.5 Anthropic [2025], Doubao-Seed-1.8 ByteDance Seed [2025], MiMo-V2-Flash Team et al. [2026], Qwen3-VL-235B-A22B-Instruct Bai et al. [2025], and Qwen3-VL-8B-Instruct Bai et al. [2025].
  • Tool-Augmented MLLMs. Additionally, we evaluate tool-augmented model services accessed through their official web platforms, with the maximum reasoning mode enabled to elicit their full capabilities. Concretely, we evaluate the following models: GPT-5.2-Thinking OpenAI [2026], Gemini-3-Pro-Preview Google [2026], and Claude-Sonnet-4.5-Thinking Anthropic [2026].
@@ 4.2 Main Results
(5) Competitive Performance of Open-Source Models. While proprietary models (e.g., GPT-5.2 OpenAI [2025b]) remain the top performers, high-performance open-source models are rapidly closing the gap. Notably, Doubao-Seed-1.8 ByteDance Seed [2025] achieves $33.67\%$ SR when equipped with OmniSeeker. This demonstrates that high-quality open-source models possess strong capacity for complex reasoning and provide a promising path toward developing cost-effective, high-performance web browsing agents.

### v2

@@ 1 Introduction
Multimodal large language models [OpenAI, 2025; Pichai et al., 2025; Bai et al., 2025; Meta, 2025; Li et al., 2024a; Shi et al., 2025a] have demonstrated substantial performance gains across complex tasks. By integrating linguistic comprehension, visual perception, and tool-use capabilities, these models are increasingly evolving into autonomous agents capable of independent exploration and decision-making. Consequently, an increasing body of research is exploring how MLLMs can leverage external search and browsing tools to address multimodal deepsearch challenges in open-world environments [OpenAI, 2025; Google, 2024; Wu et al., 2025a; Geng et al., 2025; Huang et al., 2026].
Despite the rapid evolution of model capabilities, benchmarks for multimodal browsing and deep search remain noticeably underdeveloped. Existing studies [Jiang et al., 2024; Geng et al., 2025; Li et al., 2025a; Tao et al., 2025a] frequently exhibit shortcomings in task complexity, information searchability, and evaluation dimensions, hindering fair, holistic, and reproducible assessments of multimodal browsing agents. Existing methods still exhibit certain limitations: i) Insufficient Task Complexity. Early benchmarks [Cheng et al., 2025; Geng et al., 2025] are predominantly confined to shallow retrieval within two hops, with visual information concentrated in the initial stage. Consequently, they fail to reflect the intricacies of real-world, deep multimodal search scenarios. ii) Inaccessibility of Key Information. The core evidence in subsequent benchmarks [Fu et al., 2025a; Li et al., 2025a; Tao et al., 2025a] is often derived from sources that are not publicly searchable by tools, such as videos or proprietary documents, which undermines the reproducibility and fairness. iii) Narrow Evaluation Dimensions. Existing studies [Jiang et al., 2024; Li et al., 2024b] primarily focus on the accuracy of the final answer but lack a systematic characterization of the reasoning process. This makes it challenging to diagnose specific failure modes or define the capability boundaries of the models.
@@ 2.1 Multimodal Large Language Models
Multimodal large language models [OpenAI, 2025; Pichai et al., 2025; ByteDance Seed, 2025; Bai et al., 2025] have demonstrated remarkable proficiency across a diverse spectrum of tasks, such as VQA [Fu et al., 2025b; Cheng et al., 2025; Zhang et al., 2025; Wu et al., 2025a], grounding [Kazemzadeh et al., 2014], OCR [Masry et al., 2022; Shi et al., 2025b; Mathew et al., 2021; Shi et al., 2025c], and multimodal reasoning [Lu et al., 2023; Wang et al., 2025, 2024]. Nevertheless, MLLMs inherently struggle with the real-time acquisition of up-to-date information, posing substantial hurdles when addressing knowledge-intensive or information-retrieval queries. Consequently, contemporary research has pivoted toward tool-augmented frameworks to empower MLLMs as autonomous agents, capable of dynamically retrieving and incorporating external knowledge.
@@ 2.2 Tool-Enhanced Browsing Agents
Driven by the escalating tool-calling proficiency of LLMs/MLLMs [Guo et al., 2025; Yang et al., 2025; OpenAI, 2025], tool-enhanced browsing agents have emerged as a pivotal research frontier. To enable precise retrieval and reasoning in dynamic web environments, recent studies advocate leveraging supervised fine-tuning and reinforcement learning to enhance agents’ reasoning and decision-making capabilities [Jin et al., 2025; Li et al., 2025b; Wu et al., 2025b; Tao et al., 2025b; Song et al., 2025; Zheng et al., 2025]. This paradigm, initially validated in textual agents, has been rapidly extended to the multimodal domain [Wu et al., 2025a; Mei et al., 2025; Hong et al., 2025; Huang et al., 2026]. This evolution has significantly expanded the search depth and adaptive boundaries of agents when navigating complex tasks.
@@ 2.3 Multimodal Browsing Benchmarks
Traditional multimodal browsing benchmarks typically decouple visual understanding from text retrieval and focus on simple two-hop retrieval tasks [Chen et al., 2023; Mensink et al., 2023; Cheng et al., 2025; Geng et al., 2025]. As visual agents advance, performance on such tasks has largely saturated. BrowseComp [Wei et al., 2025] evaluates text-only agents in open-world settings by requiring large-scale web navigation, offering valuable guidance for multimodal task construction. Inspired by this paradigm, recent benchmarks such as MM-BrowseComp [Li et al., 2025a] and MMSearch-Plus [Tao et al., 2025a] incorporate multi-hop designs and fine-grained visual reasoning to enhance reasoning depth. However, existing benchmarks still suffer from key limitations: critical information often resides in videos or non-searchable documents, tool support is insufficient, and evaluation primarily measures final-answer correctness while overlooking the quality of reasoning. To bridge this gap, we propose BrowseComp-V3, ensuring all critical information comes from publicly accessible resources during task design. We also introduce the Process Score metric to evaluate multimodal browsing agents comprehensively.
@@ 4.1 Experimental Setup
  • Tool-Free MLLMs. We benchmark multiple SOTA MLLMs in a tool-free setting, where models must generate answers directly without access to external tools or search capabilities. Specifically, we evaluate the following models: GPT-5.2 [OpenAI, 2025], o4-mini [OpenAI, 2025], GPT-4o [Hurst et al., 2024], Gemini-3-Flash-Preview [Google, 2025], Claude-Sonnet-4.5 [Anthropic, 2025], Doubao-Seed-1.8 [ByteDance Seed, 2025], MiMo-V2-Flash [Team et al., 2026], Qwen3-VL-235B-A22B-Instruct [Bai et al., 2025], and Qwen3-VL-8B-Instruct [Bai et al., 2025].
  • Tool-Augmented MLLMs. Additionally, we evaluate tool-augmented model services accessed through their official web platforms, with the maximum reasoning mode enabled to elicit their full capabilities. Concretely, we evaluate the following models: GPT-5.2-Thinking [OpenAI, 2026], Gemini-3-Pro-Preview [Google, 2026], and Claude-Sonnet-4.5-Thinking [Anthropic, 2026].
@@ 4.2 Main Results
(5) Competitive Performance of Open-Source Models. While proprietary models (e.g., GPT-5.2 [OpenAI, 2025]) remain the top performers, high-performance open-source models are rapidly closing the gap. Notably, Doubao-Seed-1.8 [ByteDance Seed, 2025] achieves $33.67\%$ SR when equipped with OmniSeeker. This demonstrates that high-quality open-source models possess strong capacity for complex reasoning and provide a promising path toward developing cost-effective, high-performance web browsing agents.